2/3x + 4/5= 9/10 I don't understand it at all...please help!
Answer by rfer(16322)

2/3x+4/5=9/10
find common denominator
20/30x+25/30=27/30
subtract 25/30
20/30x=2/30
divide by 20/30
to do that we invert and multiply
x=(2/30)/(20/30)
x=(2/30)(30/20)
x=60/600
reduce fraction
x=1/10